Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
6x + 2y = 2
3x + 5y = 5

Answer by Alan3354(69443)

If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
6x + 2y = 2 eqn 1
3x + 5y = 5 eqn 2
--------------
6x +10y = 10 eqn 2 times 2
6x + 2y = 2 eqn 1
0x + 8y = 8 Subtract eqn 1
y = 1
--------
3x + 5*1 = 5
x = 0
(x,y) = (0,1)